OpenStudy (abbycross167):

Can someone please help me? I really need help Find the point-slope form of the equation that goes through these two points. (1, 20) and (8, 4.5)

OpenStudy (4everaddicted2anime):

The formula is y-(y coordinate)=m(x-(x coordinate)) You need to find slope first: Points: (1,20)(8,4.5) From these points you need to plug them into the slope formula which is your 2nd y coordinate minus the 1st y cooridinate all over the 2nd x coordinate minus the 1st coordinate. (4.5-20)/(8-1) this comes out to -5.5 over 7. Now you just plug the slope and one of the coordinate pairs into the formula. Once you're done you get y-20=-55 over 7(x-1).

OpenStudy (abbycross167):

Thank you soo much!!